OpenStudy (anonymous):

How to solve b^2=-6-7b

OpenStudy (ryanchick):

flip the problem around

OpenStudy (anonymous):

How do I do that

OpenStudy (ryanchick):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : b^2-(-6-7*b)=0

OpenStudy (anonymous):

Is that they answer

OpenStudy (ryanchick):

Factoring b2+7b+6 The first term is, b2 its coefficient is 1 . The middle term is, +7b its coefficient is 7 . The last term, "the constant", is +6 Step-1 : Multiply the coefficient of the first term by the constant 1 • 6 = 6 Step-2 : Find two factors of 6 whose sum equals the coefficient of the middle term, which is 7 . -6 + -1 = -7 -3 + -2 = -5 -2 + -3 = -5 -1 + -6 = -7 1 + 6 = 7 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, 1 and 6 b2 + 1b + 6b + 6 Step-4 : Add up the first 2 terms, pulling out like factors : b • (b+1) Add up the last 2 terms, pulling out common factors : 6 • (b+1) Step-5 : Add up the four terms of step 4 : (b+6) • (b+1) Which is the desired factorization

OpenStudy (mathmale):

b^2=-6-7b is an example of a very important category of math function: It's a quadratic. What does a quadratic equation look like when written in standard form? b^2=-6-7b needs to be re-written in descending powers of b.

OpenStudy (anonymous):

I'm lost

OpenStudy (ryanchick):

here is the answer b^2-(-6-7*b)=0 do you understand

OpenStudy (anonymous):

Thx for the but no I don't I sux at math

OpenStudy (anonymous):

Suck

OpenStudy (ryanchick):

all you practically do is flip it around

OpenStudy (anonymous):

Wat about this one b^2-9=0

OpenStudy (ryanchick):

b^(2)-(9)=0 so just rearrange and your answer should be b=3

OpenStudy (ryanchick):

is that good